[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO Status Report for December , 1995

   Compton Gamma Ray Observatory Status Report #188
             Friday December 8, 1995

      Questions or comments can be sent to 
          Chris Shrader at the CGRO-SSC.
          Phone:  301/286-8434
          e-mail: shrader@grossc.gsfc.nasa.gov
            Guest Investigator News

The Cycle-6 NRA release is imminent! The document
and its appendices are currently going through
approval cycle at Goddard, after which it will be
sent for final review by the Program Scientist. A
December 22, 1995 release, with a March 22, 1996
proposal due date is now the most likely
scenario. As soon it is formerly approved, the
NRA will be mailed out, and made available
electronically (through the GRO-SSC WWW site at
or GRONEWS, grossc.gsfc.nasa.gov, gronews).

The current Target of Opportunity re-orientation
of the space craft (refer to the BATSE report
below) was a yaw-only maneuver, and thus does not
effect any COMPTEL or EGRET integrations. 

The CGRO Exhibit at the Natinal Air and Space
Museum is scheduled to open on JANUARY 9, 1996.
It consists of a 6-minute video on CGRO, and
still visuals and text describing scientific
highlights of the mission. It will be on display
through the summer of 1996 in the "Milestones of
Flight" hall of the museum.


EGRET operations were normal this monthly period. 
Delivery of the final phase 3 data to the GRO SSC
has been completed on schedule, and delivery of
the phase 4 preliminary data to the GRO SSC is
also on schedule. Interaction with guest
investigators continues at a good level.

During the period from October 31, 1995 to
November 28, 1995, EGRET was turned off to
conserve the remaining spark chamber gas while
observations were being made which were of
primary interest to those using the other
instruments on GRO.  On November 28, EGRET was
activated and is currently pointing in the
direction defined by (l=77.44 degrees, b=-38.58
degrees). There was a rotation about the z axis
on December 6 for a target of opportunity, but
that did not effect the collection of EGRET data
for this region.  Starting on December 12, EGRET
will be observing the region centered on (l=5.50
degrees, b=-0.17 degrees).

Scientific papers on bright supernovae remnants
and the pulsar B9656+14 have been accepted by the
Astrophysical Journal, and an article on the
evolution of gamma-ray loud active galactic
nuclei has appeared in the Astrophysical Journal.


The COMPTEL operations group reports that an ADC
in COMPTEL's Remote COMPTEL Interface Unit-A
(RIU-A), external to the instrument itself on the
spacecraft platform, failed on 24 November.  This
ADC provides a monitor to the POCC at NASA/GSFC
of COMPTEL detector assembly temperatures and
power supply voltages.  A similar failure
occurred in March 1994 for COMPTEL's RIU-B. 
Independent COMPTEL instrument telemetry
continues to provide the relevant experiment
state-of-health information without problem. 
There has been no interruption to the flow or
quality of COMPTEL's scientific data.  Operations
procedures have been adjusted to compensate for
this loss of monitoring capability.

A number of additional gamma-ray bursts have
occurred within the field of view of COMPTEL
since the last report (GRBs 951117, 951119,
951128, 951202, and 951208).  Once again, none of
these was detected at MeV energies by COMPTEL. 
The last burst detected within the field of view
of COMPTEL occurred in May 1995, establishing a
new record (of sorts) for the longest period
without a burst detection by the telescope. Since
May 1995, the COMPTEL burst group has responded
to a total of 25 BATSE/BACODINE triggers with
locations falling within the COMPTEL field of
view; typical response times are now on the order
of 15 minutes after a BATSE trigger.  Procedures
are currently being implemented to further reduce
this response time to approximately 3 to 5
minutes following receipt of a burst trigger
message.  A new release of COMPTEL data to the
CGRO public archive is scheduled for the near
future. This release will include both low- and
high-level data products, in anticipation of the
CGRO Cycle 6 proposal cycle.


OSSE operations are normal.  The instrument is
working as designed, with all subsystems in
complete and full operation. The slewing response
to BATSE burst triggers has been enabled since 30

OSSE 2 minute spectral data products are being
reprocessed to fix miscalculations in ODS data
(data stored on-board during TDRSS realtime
outages) beginning with viewing period 211.  The
corrections affect some  medium and high range
livetimes and some spectral channel counts during 
approximately 1% of a typical day of data. 
Recent observations are listed in the following

View period      Dates        Target (owner)
    503         14-21 Nov     N Ecliptic Pole Survey          
                  		(PI team)
                              M 31 (not assigned)
                              NGC 5506 (not assigned)
    504         21-28 Nov     N Ecliptic Pole survey              
                  		(PI team) 
                              M 31 (not assigned)
    507      28 Nov - 8 Dec   CTA 102 (not assigned)
                              GRS 1915+105  (PI team)
                              NGC 5506  (not assigned)
                              (A slight yaw was                  
          		      approved for this viewing 
                              period to give OSSE a           	
		              better view of the        
                              recurring x-ray transient
                              source GRS 1915+105).

Last month it was reported that viewing period
502 targets included Geminga,  NGC 3227, NGC
4388, M 87, and NGC 1068.  NGC 1068 was not
observed during this viewing period.

Data through viewing period 338 have been
delivered to the Compton GRO Science  Support
Center archive.


The following notices appeared in IAU Circular

GRS 1915+105                

B. A. Harmon, W. S. Paciesas, S. N. Zhang and K.
J. Deal,  NASA/Marshall Space Flight Center,
report, for the Compton  Observatory BATSE Team:
"Hard x-ray flux (20-100 keV) from GRS  1915+105
has increased over the past month to a level
comparable to the intense outbursts seen in 1992
and 1994 (IAUC 5590, 5959). The flux reached 300
mCrab around Oct. 15 and has remained near this
level (within 50 percent) as of Nov. 20. Less
intense  outbursts reaching about 200 mCrab have
also been observed since June 1995. These are
nonperiodic but are of similar intensity  and
each about 15-20 days in duration. The source
continues to be variable on a daily basis. The
hard x-ray activity correlates well on long
timescales (days to weeks) with the radio flux,
but no strong signature is seen in x-rays that
corresponds to the shorter- timescale (hours to
days) radio flares reported below."  
4U 0115+634    

M. H. Finger, M. Scott and K. Hagedon,
Universities Space Research Association; R. B.
Wilson and C. A. Wilson, NASA/Marshall Space
Flight Center; M. Stollberg and W. S. Paciesas,
University of Alabama, Huntsville; and T. A.
Prince and B. Vaughan, California Institute of
Technology, report for the BATSE/Compton GRO
Observatory pulsar team: "Pulsed hard x-ray flux
is currently being detected from the transient
x-ray pulsar 4U 0115+634. The source was detected
on Nov. 18, when the rms pulsed flux (20-50 keV)
was approximately 15 mCrab. This is the first
detection by BATSE of this source since the
outburst that occurred during 1994 May 8-June 27
(IAUC 5990, 5999). On 1995 Nov. 21 the rms pulsed
flux (20-50 keV) was approximately 70 mCrab. The
pulse frequencies observed from Nov. 19 to 21 are
consistent with a neutron- star spin rate of
0.2766635 +/- 0.0000003 Hz at epoch JD 2450042
and a spin-up rate of 1.4 +/- 0.3 x 10E-11 Hz/s,
using the binary orbital ephemeris of Cominsky et
al. (1994, 2nd Compton Symposium, AIP Proc. 304,
Currently an unusual source near the galactic
center region is being detected. The source has a
soft spectrum and is not detected above 50 keV.
Efforts are underway to determine the source
location.  A spacecraft re-orientation was 
requested and granted in order to improve BATSE's
source location sensitivity.

During the past 30 days (TJD 10028 - 100058) the
following pulsed sources have been detected by
the BATSE pulsed source monitor: Her X-1, 4U
0115+634,  Cen X-3, 4U 1626-67, OAO 1657-415, GX
1+4, Vela X-1, and GX 301-2.

As of December 6th BATSE has detected 1425
gamma-ray bursts out of a total of 3828 on-board
triggers in 1688 days of operation. There have
been 766 triggers due to solar flares, 10 due to
SGR events, and 51 due to terrestrial gamma-ray